julio is 36 years old, and james is 11 years old. IN how many years woill Julios age be twice james age?
Found 2 solutions by reynard2007, Alan3354:
Answer by reynard2007(52)


let x be the number of years to pass...
in x years, both of them will x years older, so:
1)Julio will be 36+x years old and
2) James will be 11+x years old


36+x = 2(11+x)
36+x = 22+2x
36-22 = 2x - x
x = 14

Answer by Alan3354(69443)

julio is 36 years old, and james is 11 years old. IN how many years woill Julios age be twice james age?
--------------
The age difference is 25 years.
--> when they're 25 and 50